OpenStudy (nora):

f(x)=x sin x(x^2) Find f^(2011) (0).

OpenStudy (amistre64):

2011/360=?

OpenStudy (amistre64):

5.58611 degrees; but is 2011 in radians or degrees for that?

OpenStudy (nora):

I believe you have to work this using Taylor series

OpenStudy (amistre64):

oh.... then you prolly want someone else to help you ; like the cow :)

OpenStudy (nora):

rude.

OpenStudy (amistre64):

no..... rude would be me doing something absurd and uncalled for; this is just me not being good with taylor series and suggesting you find someone smarter than me...

OpenStudy (amistre64):

like...the cow here :)

OpenStudy (nora):

oh ok ok i mis understood u then . thanks for attempting

OpenStudy (amistre64):

sorry i couldnt help out :) good luck with it tho ;)

OpenStudy (anonymous):

Just for clarification, is it \[x \sin (x*x ^{2})\]

OpenStudy (nora):

no its (x)(sin(x^2)

OpenStudy (nora):

)

OpenStudy (anonymous):

so if you taylor expand this, you will get\[x(x^2 - (x^2)^3/3! - (x^2)^5/5! + ...)\]

OpenStudy (anonymous):

and if you plug in 0 for this you get 0 .-.

OpenStudy (nora):

did u use a formula

OpenStudy (anonymous):

well, I just used the taylor expansion of sin x and then I plugged in x^2 for that x

OpenStudy (nora):

then multiplied the series by x right??

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

and because you can always factor out an x term, when you plug in x, you'll get 0, or at least this is my idea of how it works

OpenStudy (anonymous):

"no its (x)(sin(x^2)" f(x)= x*sin(x^2) f(0) = 0. At the origin both x x^2 are zero. Thus the product of x and Sin[x^2] is zero. Refer to the attached plot, Nora_1.pdf .

OpenStudy (anonymous):

Robtobey, I think the problem isn't taking the function to the 2011 power, but rather taking the 2011 derivative